# Semi-parametric Bernstein-von Mises in Linear Inverse Problems

### v2

@@ 3.2 Location Recovery in Semi-blind Deconvolution
Let $H=L^{2}[0,1]$ and let $\Theta\subset(-1/2,1/2)$ be compact. Let $\langle f_{0},e_{1}\rangle\neq 0$, $g_{1}\neq 0$ and $\sup_{k}k^{3/2+\delta}|g_{k}|<\infty$, for some $\delta>0$. Then the solution operator $K_{\theta}$ given in (25) satisfies Assumption 2.2 relative to $S=S^{\eta}$ for $\eta\leq 1$. Furthermore, (8) is satisfied for $\eta>1/2$, the map $\theta\mapsto\|K_{\theta}f\|$ is constant for every $f$, $\dot{K}_{\theta_{0}}f_{0}\neq 0$ and $(K_{\theta}-K_{\theta_{0}})f_{0}\neq 0$ for $\theta\neq\theta_{0}$.
@@
(i). By 1-periodicity $\|K_{\theta}f\|=\|g*f(\cdot-\theta)\|=\|f*g\|$, which is independent of $\theta$. Here $\|f*g\|\leq\|f\|\,\|g\|$, by the Cauchy-Schwarz inequality (and periodicity), whence $\|K_{\theta}\|=\sup_{\|f\|\leq 1}\|f*g\|\leq\|g\|<\infty$. Similarly $\|\dot{K}_{\theta}\|\leq\|g^{\prime}\|<\infty$. (ii). Choose $1/2<\eta\leq 1/2+\delta$, so that the sequence $|g_{k}|k^{\eta+1}$ is bounded. Then
@@
by the inequality $|e^{ix}-e^{iy}|\leq|x-y|$, for every $x,y\in\mathbb{R}$. (iii). By the exact formulas
@@ 4.2 $p$-exponential priors
If the Fourier coefficients of $g$ are of polynomial order $|k|^{-q}$, then we require $q>3/2$ for Proposition 3.2 to be applicable. (We changed notation, denoting the order of the Fourier coefficients now by $q$ instead of $p$, as $p$ refers to the prior here.) We assume that $f_{0}\in S^{\beta}$ and then have that $\gamma_{\theta_{0},f_{0}}\in S^{\beta+2q-1}$, in view of (27). Lemma 5.13 in [1] then gives
@@
Whenever $\beta,q$ are not small, this limit dominates $A(\beta,q,2)$. For example, if we only use that $q\geq 2$, then this happens if $\beta>(\sqrt{11}-2)/2$.
@@ 6 Proofs
$\qquad+\sqrt{n}(\theta-\theta_{0})\langle\dot{W}^{(2)},(K_{\theta_{0}}-K_{\theta})\gamma_{n}\rangle$
@@
$\qquad+\frac{n}{2}(\theta-\theta_{0})^{2}\|\dot{K}_{\theta_{0}}f\|^{2}-\frac{n}{2}\|(K_{\theta}-K_{\theta_{0}})f\|^{2}$
$\qquad+\frac{n}{2}(\theta-\theta_{0})^{2}\bigl{(}\|\dot{K}_{\theta_{0}}f_{0}\|^{2}-\|\dot{K}_{\theta_{0}}f\|^{2}\bigr{)}.$
@@
Finally we prove (33), where we consider the six terms defining $R_{n}(\theta,f)$ separately.
@@
The second term of (33) is a centered Gaussian variable with variance $n|\theta-\theta_{0}|^{2}\,\|(K_{\theta}-K_{\theta_{0}})\gamma_{n}\|^{2}\lesssim n|\theta-\theta_{0}|^{4}\|\gamma_{n}\|^{2}$, by (ii) of Assumption 2.2. Thus the second term is (easily) $o_{P}\bigl{(}1+n(\theta-\theta_{0})^{2})$.
In the third term we bound $n|\theta-\theta_{0}|/(1+n|\theta-\theta_{0}|^{2})$ by $\sqrt{n}$, and thus reduce it to assumption (10).
By the Cauchy-Schwartz inequality and item (iii) of Assumption 2.2, the fourth term is bounded above by a multiple of $n(\theta-\theta_{0})^{2}\|f\|\,\|f-f_{0}\|$, which is $o(1+n(\theta-\theta_{0})^{2})$, uniformly in $f\in H_{n}$.
Similarly, by item (ii) of Assumption 2.2, the fifth term is bounded above by a multiple of $n|\theta-\theta_{0}|^{3}\|f\|$, which is $o(1+n(\theta-\theta_{0})^{2})$ uniformly on $\Theta_{n}\times H_{n}$.
Finally the sixth term of (33) is bounded by $n(\theta-\theta_{0})^{2}\|\dot{K}_{\theta_{0}}(f-f_{0})\|\,\|\dot{K}_{\theta_{0}}(f+f_{0})\|$, which is $o(1+n(\theta-\theta_{0})^{2})$ by item (i) of Assumption 2.2.
@@
[Acknowledgments]
The authors would like to thank Frank van der Meulen for his helpful assistance and comments.

### v3

# Semi-parametric Bernstein-von Mises Theorem in Linear Inverse Problems

@@ Abstract
62F15,
62E20,
Bayesian statistics,
inverse problems,
semi-parametric statistical models,
Bernstein-von Mises theorem,
heat equation,
deconvolution,
keywords:
[class=MSC]
keywords:
@@ 3.2 Location Recovery in Semi-blind Deconvolution
Let $H=L^{2}[0,1]$ and let $\Theta\subset(-1/2,1/2)$ be compact. Let $\langle f_{0},e_{1}\rangle\neq 0$, $g_{1}\neq 0$ and $\sup_{k}k^{5/2+\delta}|g_{k}|<\infty$, for some $\delta>0$. Then the solution operator $K_{\theta}$ given in (25) satisfies Assumption 2.2 relative to $S=S^{\eta}$ for $\eta\leq 1/2+\delta$. Furthermore, (8) is satisfied for $\eta>1/2$, the map $\theta\mapsto\|K_{\theta}f\|$ is constant for every $f$, $\dot{K}_{\theta_{0}}f_{0}\neq 0$ and $(K_{\theta}-K_{\theta_{0}})f_{0}\neq 0$ for $\theta\neq\theta_{0}$.
@@
(i). By 1-periodicity $\|K_{\theta}f\|=\|g*f(\cdot-\theta)\|=\|f*g\|$, which is independent of $\theta$. Here $\|f*g\|\leq\|f\|\,\|g\|$, by the Cauchy-Schwarz inequality (and periodicity), whence $\|K_{\theta}\|=\sup_{\|f\|\leq 1}\|f*g\|\leq\|g\|<\infty$. Similarly $\|\dot{K}_{\theta}\|\leq\|g^{\prime}\|<\infty$. (ii). Choose $1/2<\eta\leq 1/2+\delta$. The sequence $|g_{k}|k^{\eta+1}$ is thus bounded. Then
@@
by the inequality $|e^{ix}-e^{iy}|\leq|x-y|$, for every $x,y\in\mathbb{R}$. (iii). As $\eta\leq 1/2+\delta$, the sequence $|g_{k}|k^{\eta+2}$ is bounded. By the exact formulas
@@ 4.2 $p$-exponential priors
If the Fourier coefficients of $g$ are of polynomial order $|k|^{-q}$, then we require $q>5/2$ for Proposition 3.2 to be applicable. (We changed notation, denoting the order of the Fourier coefficients now by $q$ instead of $p$, as $p$ refers to the prior here.) We assume that $f_{0}\in S^{\beta}$ and then have that $\gamma_{\theta_{0},f_{0}}\in S^{\beta+2q-1}$, in view of (27). Lemma 5.13 in [1] then gives
@@
Whenever $\beta,q$ are not small, this limit dominates $A(\beta,q,2)$. For example, if we only use that $q\geq 3$, then this happens if $\beta>\sqrt{27}/2-2$.
@@ 6 Proofs
$\qquad+\frac{n}{2}(\theta-\theta_{0})^{2}\|\dot{K}_{\theta_{0}}f\|^{2}-\frac{n}{2}\|(K_{\theta}-K_{\theta_{0}})f\|^{2}.$
@@
Finally we prove (33), where we consider the four terms defining $R_{n}(\theta,f)$ separately.
@@
For the second term, we bound $n|\theta-\theta_{0}|/(1+n|\theta-\theta_{0}|^{2})$ by $\sqrt{n}$, and thus reduce it to assumption (10).
By the Cauchy-Schwartz inequality and item (iii) of Assumption 2.2, the third term is bounded above by a multiple of $n(\theta-\theta_{0})^{2}\|f\|\,\|f-f_{0}\|$, which is $o(1+n(\theta-\theta_{0})^{2})$, uniformly in $f\in H_{n}$.
Finally, by item (ii) of Assumption 2.2, the fourth term is bounded above by a multiple of $n|\theta-\theta_{0}|^{3}\|f\|$, which is $o(1+n(\theta-\theta_{0})^{2})$ uniformly on $\Theta_{n}\times H_{n}$.
@@
[Acknowledgements]
The authors would like to thank Frank van der Meulen for his helpful assistance and comments, as well as the two anonymous referees and the Associate Editor for their constructive comments that improved the quality of this paper.
{funding}
This research is partly supported by NWO grant 613.009.134 awarded to J.H. van Zanten and the NWO Spinoza prize awarded to A.W. van der Vaart by the Netherlands Organisation for Scientific Research (NWO).